# references.bib AIJIM: A Theoretical Model for Real-Time, Crowdsourced Environmental Journalism with AI

### v3

@@ Abstract
Environmental journalism is vital for raising awareness of ecological crises and driving evidence-based policy, yet traditional methods falter under delays, inaccuracies, and scalability limits, especially in under-monitored regions critical to the United Nations Sustainable Development Goals [IPCC2023, Robbins2023, Dai2024].
To bridge these gaps, this paper introduces the Artificial Intelligence Journalism Integration Model (AIJIM), a conceptual and transferable model for structuring AI-enhanced workflows in environmental journalism. AIJIM combines real-time hazard detection, dual-level validation (visual and textual), and AI-driven reporting, independent of specific technologies or implementation platforms.
Validated through a pilot study, AIJIM significantly improved the speed and accuracy of environmental hazard reporting, outperforming traditional methods. Furthermore, the model directly addresses key ethical, regulatory, and scalability challenges, ensuring AI accountability through Explainable AI (XAI), GDPR-compliant data governance, and active public participation.
By explicitly incorporating both crowdsourced visual validation and human review of AI-generated textual reports, AIJIM provides a uniquely transparent and adaptable solution, setting a new benchmark for AI-supported environmental journalism and fostering informed global decision-making across diverse socio-political landscapes.
Index Terms:
Artificial Intelligence, Environmental Journalism, Real-Time Reporting, Vision Transformers, Image Recognition, Crowdsourced Validation, GPT-4, Automated News Generation, GIS Integration, Data Privacy Compliance, Explainable AI (XAI), AI Ethics, Sustainable DevelopmentI Introduction
@@
Environmental journalism plays a crucial role in raising awareness and shaping policy regarding ecological crises such as climate change, pollution, and biodiversity loss, influencing public discourse and evidence-based decision-making [IPCC2023, Robbins2023]. However, traditional reporting methods face significant limitations, particularly delays in data acquisition, costly verification processes, and limited scalability. These constraints are especially problematic in under-monitored regions vital to global sustainability, delaying response to urgent environmental threats such as oil spills or illegal deforestation [Reith2025].
@@
Artificial Intelligence (AI), including Natural Language Processing (NLP), Computer Vision (CV), and Machine Learning (ML), has already transformed journalism by enhancing automation, data processing, and analytical capabilities. Nevertheless, existing AI-driven journalistic frameworks are constrained by limited adaptability and insufficient real-time processing capabilities, particularly in rapidly evolving environmental contexts.
To address these critical challenges, this paper introduces the Artificial Intelligence Journalism Integration Model (AIJIM), a modular and transferable theoretical model designed to guide the integration of AI into real-time environmental journalism. AIJIM combines advanced AI analysis with community-driven validation to ensure high accuracy, rapid response, and ethical accountability. Unlike traditional AI journalism approaches, AIJIM processes real-time citizen-generated imagery and adopts a technology-agnostic structure, enabling seamless adaptation to future AI advancements.
The effectiveness of AIJIM was demonstrated through a pilot study conducted in Mallorca, validating its practical applicability and responsiveness to real-world environmental hazards. By emphasizing adaptability and community engagement, AIJIM uniquely addresses existing gaps in Data-Driven Journalism (DDJ), Computational Journalism, and AI Fact-Checking models, which largely focus on structured, retrospective data or textual claims.
This paper evaluates AIJIM’s methodological and practical contributions, highlighting its transformative potential for enhancing responsiveness, accuracy, and credibility in environmental journalism. Section III provides a comprehensive overview of AIJIM, outlining its theoretical foundation and conceptual components.
II Background and Related Work
@@
Existing research underscores the increasing integration of Artificial Intelligence (AI) in journalism, yet significant limitations persist. Traditional AI approaches, such as Data-Driven Journalism (DDJ), depend primarily on structured, retrospective datasets, rendering them ineffective for real-time environmental reporting [Ivancsics2019, Tong2022]. Similarly, Computational Journalism, which employs machine learning for complex pattern recognition, struggles with the dynamic and unstructured nature of live environmental data [Mudzanani2025, Hidayati2025]. Additionally, AI Fact-Checking models, focusing on textual claim verification, lack capability for real-time visual or spatial hazard detection [Wolfe2024, Cazzamatta2025].
@@
Building on these insights, it is crucial to evaluate how existing AI-driven journalism frameworks address these challenges. This section identifies their methodological constraints and positions the Artificial Intelligence Journalism Integration Model (AIJIM) as a novel approach to integrating real-time automation, adaptability, and scalable journalistic workflows.
II-A AI in Journalism
@@
AI technologies have significantly advanced journalism by enabling automation, rapid analysis, and streamlined verification processes. Natural Language Processing (NLP) has enhanced automated news generation from structured datasets, significantly improving journalistic efficiency [OpenAI2023, Tong2022]. Additionally, developments in Computer Vision (CV) have facilitated rapid detection and classification of visual data, surpassing traditional manual methods in both speed and accuracy [Tiltack2025, Floridi2021]. Despite these advancements, current AI-driven methods often rely heavily on periodic updates from structured sources, such as satellite imagery, limiting their effectiveness in dynamic and rapidly evolving environmental scenarios. This gap emphasizes the critical need for AI solutions like AIJIM, which can efficiently handle dynamic and real-time data streams.
II-B Challenges in Environmental Journalism
@@
Environmental journalism frequently encounters significant practical challenges, especially concerning the timely acquisition and verification of reliable data. This issue is particularly pronounced in remote or under-monitored regions, complicating rapid responses to urgent environmental threats [Robbins2023]. Conventional journalism methods, characterized by manual data collection and expert-driven verification, are resource-intensive, subject to inaccuracies, and hinder scalability and timely intervention [Ivancsics2019, deLimaSantos2022]. Consequently, there is an urgent demand for innovative AI-driven solutions to improve the precision, timeliness, and overall credibility of environmental reporting.
II-C Limitations of Existing AI Models
@@
Most AI models for environmental monitoring rely on satellite-based imagery and long-term trend analysis, making them effective for large-scale environmental assessments but insufficient for detecting localized, high-frequency hazards like illegal dumping or micro-scale pollution [Ivancsics2019, Tong2022]. Additionally, existing AI solutions often fail to adequately adjust for contextual variability, leading to misclassifications in underrepresented geographic regions [deLimaSantos2022]. Regulatory constraints, including GDPR compliance and data privacy risks, further limit the scalability of these models [Floridi2021].
@@
II-D Comparison of AIJIM with Existing Journalism Models
@@
Although several AI-based journalism models have been proposed, none fully meet the specific demands of real-time environmental crisis reporting. Data-Driven Journalism (DDJ), focused on historical structured datasets, lacks immediate adaptability [Tong2022]. Similarly, Computational Journalism, while proficient at analyzing extensive datasets, struggles with dynamic, real-time environmental scenarios [Sallami2024]. Meanwhile, AI Fact-Checking excels primarily in textual verification but falls short in processing real-time visual and spatial data critical for environmental monitoring [Cazzamatta2025].
@@
AIJIM addresses these limitations by uniquely integrating real-time hazard detection, adaptive community validation, and automated reporting into a coherent, modular framework. This combination ensures scalability, adaptability, and ethical AI governance, bridging the gap between automation and human oversight in environmental journalism. A detailed methodological comparison is presented in Section III.
@@
Importantly, unlike purely technical frameworks or platform-specific toolkits, AIJIM is deliberately conceptualized as a theoretical model. Its goal is to offer a transferable structure that enables systematic analysis and design of AI-supported environmental journalism workflows. By integrating technological, ethical, and participatory dimensions, AIJIM contributes to ongoing efforts in theory-building at the intersection of journalism studies, human-centered AI, and digital environmental communication.
III The Artificial Intelligence Journalism Integration Model (AIJIM)
@@
The Artificial Intelligence Journalism Integration Model (AIJIM) advances beyond conventional AI-driven journalism approaches by integrating real-time hazard detection, automated validation, and AI-generated reporting into a scalable, modular, and adaptable model structure. Unlike traditional models, AIJIM is designed for dynamic, real-time environmental journalism, allowing for flexible integration of AI technologies tailored to specific reporting needs.
AIJIM’s architecture supports multiple AI components to ensure technological adaptability. In this study, AIJIM leverages Vision Transformer-based DETR models for high-precision environmental hazard detection, providing both high accuracy and low latency [Tiltack2025]. However, the model is not limited to a single AI configuration and supports alternative architectures based on computational efficiency and accuracy requirements.
Beyond automated detection, AIJIM enhances credibility through crowdsourced validation, ensuring robust community engagement. In the 2024 Mallorca pilot study, this approach achieved an 89.7% agreement rate with expert annotations, strengthening the accuracy and trustworthiness of AI-generated insights [Tiltack2025]. For automated reporting, AIJIM currently employs GPT-4, reducing reporting latency by and enabling rapid content generation [OpenAI2023]. Nevertheless, the model is designed to be adaptable, supporting alternative NLP models such as Claude and Gemini for diverse journalistic applications.
By enabling real-time, AI-enhanced environmental journalism, AIJIM establishes a new paradigm in which automation and human oversight converge to produce scalable, high-accuracy, and ethically responsible reporting. The model’s conceptual flexibility ensures its applicability across various journalistic environments, optimizing AI workflows for both investigative and large-scale environmental monitoring.
III-A Methodology
@@
To evaluate AIJIM’s effectiveness as a journalistic model, this study employs a Sequential Explanatory Mixed-Methods Approach [Creswell2020], combining quantitative performance data with qualitative expert perspectives:
Phase 1 (Quantitative Analysis): AIJIM’s real-time capabilities were rigorously evaluated in the 2024 Mallorca pilot study, in which 1,000 geotagged images from citizen reports enabled the detection of 50 previously undocumented waste sites—a significant advancement in AI-assisted environmental journalism [Tiltack2025, McGovern2022]. The NamicGreen platform is referenced in this study solely to demonstrate the practical applicability and real-world viability of AIJIM. It should be emphasized that the core contribution of this research is the AIJIM model itself—not the NamicGreen implementation. In this configuration, the DETR-based Vision Transformer demonstrated robust real-time detection capabilities; however, AIJIM’s modular design also supports the integration of alternative models as required.
@@
Phase 2 (Qualitative Analysis): Twelve expert interviews with professionals from journalism, AI development, and regulatory fields further informed the evaluation. A structured Qualitative Content Analysis (QCA) [Mayring2019] confirmed key thematic dimensions with a Krippendorff’s Alpha of 85%, ensuring analytical reliability.
III-B Technology Adaptability
@@
AIJIM’s architecture is designed to be adaptable to different AI models and technological environments. While this study demonstrates AIJIM’s capabilities using specific AI components, the model’s modular structure allows for the integration of alternative object detection and NLP models based on deployment constraints and accuracy requirements. Furthermore, AIJIM is compatible with diverse hardware environments, ensuring both scalability and operational flexibility across a wide range of journalistic and technical contexts. As a technology-agnostic model, AIJIM enables seamless adoption of future advancements in artificial intelligence and emerging tools without requiring structural modifications.
III-C Comparison with Existing AI-Based Journalism Models
AIJIM sets itself apart from existing AI-based journalism models—Data-Driven Journalism (DDJ), Computational Journalism, and Automated AI Fact-Checking—by redefining the scope and speed of environmental reporting. A detailed comparison is provided in Table I in Appendix A.
DDJ excels in retrospective analysis but lacks real-time responsiveness [Tong2022]. Computational Journalism offers data-driven insights yet misses dynamic hazard detection [Sallami2024]. AI Fact-Checking ensures textual accuracy but ignores visual threats [deLimaSantos2022]. AIJIM, by contrast, pioneers real-time, multi-modal journalism with active community involvement and ethical transparency, setting a new benchmark.
III-C1 Limitations of Data-Driven Journalism (DDJ)
DDJ relies on structured datasets and manual oversight, rendering it slow and unscalable for real-time environmental reporting [Stalph2023, Lewis2020]. AIJIM overcomes these by automating detection and validation.
III-C2 AIJIM vs. DDJ: A Paradigm Shift in AI Journalism
III-C3 Towards a New Standard for AI in Journalism
While AIJIM is a theoretical model designed for AI-enhanced journalism, its core principles were tested in the NamicGreen platform during the 2024 Mallorca pilot study. This proof-of-concept implementation demonstrated how AIJIM’s key components—real-time hazard detection, crowdsourced validation, and AI-driven reporting—can function in a real-world journalistic workflow, achieving a 40% latency reduction, 85.4% accuracy, and engagement of 252 validators [Verma2024, Tiltack2025]. However, AIJIM itself is not tied to this specific implementation; alternative models and platforms can be adopted to suit varying contexts. While DDJ excels in structured investigative journalism, AIJIM pioneers a scalable, automated model for real-time environmental reporting, redefining AI’s role in modern journalism. By integrating AI-driven analysis with human validation, AIJIM bridges the gap between automation and editorial judgment, ensuring that environmental journalism remains both efficient and ethically responsible. AIJIM is not just a theoretical model; its real-world applicability in NamicGreen confirms that AI-driven, real-time journalism is not only achievable but already transforming environmental reporting today.
III-C4 AIJIM vs. Other AI-Based Journalism Models: Expanding the Framework
While AIJIM significantly improves upon the Data-Driven Journalism (DDJ) paradigm, other AI-driven journalism models have also emerged. To position AIJIM within the broader AI journalism landscape, this section compares AIJIM with Computational Journalism, Automated AI Fact-Checking, and Satellite AI for Environmental Monitoring.
Computational Journalism
Computational Journalism (CJ) integrates machine learning and automation into investigative reporting, using algorithms to extract insights from large datasets [Cools2024, Nedungadi2025]. However, CJ primarily focuses on structured datasets such as financial records and political archives rather than real-time hazard detection. Unlike AIJIM, which processes live citizen-generated data, CJ relies on pre-existing databases and lacks real-time adaptability.
Strengths: Enables large-scale data-driven storytelling in fields like financial crime and political investigations [KevinAlerechi2024].
Limitations: Not optimized for real-time hazard detection; heavily dependent on structured datasets [Cools2024].
Automated AI Fact-Checking
AI Fact-Checking models, such as those deployed by Reuters, Google, and Full Fact, use Natural Language Processing (NLP) to verify claims against structured databases and news archives [Wolfe2009, Dierickx2024]. While these models improve journalistic accuracy, they do not assess real-world visual data, making them ineffective for environmental hazard detection.
Strengths: Reduces misinformation by verifying textual claims with AI-powered databases [Wolfe2009].
Limitations: Cannot validate visual hazards or environmental threats as AIJIM does [Dierickx2024].
Satellite AI for Environmental Monitoring
Satellite-based AI models (e.g., Google Earth Engine, NASA’s AI-assisted Earth Observation) have been used for large-scale environmental assessments [Dai2024, Gondwe2024]. Unlike Satellite AI, which relies on periodic remote sensing data with limited resolution, AIJIM operates on real-time, ground-level citizen-generated imagery, allowing for immediate hazard detection and validation. This significantly improves the granularity and responsiveness of environmental reporting.
Strengths: Provides large-scale, satellite-based monitoring of deforestation, urbanization, and climate change [Dai2024].
Limitations: Low resolution for small-scale hazards, limited real-time adaptability, no public engagement [Gondwe2024].
AIJIM’s Positioning Among AI Journalism Models
A detailed comparison of AIJIM with other AI-driven journalism models is provided in Table II in Appendix A.
As shown in Fig. 1, the AIJIM approach significantly outperforms other established models, especially in dimensions critical for effective environmental journalism. The radar visualization clearly highlights AIJIM’s strengths, including its capability for real-time hazard detection and active public engagement—two areas in which previous AI-driven journalism models have shown notable limitations.
Unlike Computational Journalism, AI Fact-Checking, and Satellite AI, AIJIM:
Processes real-time geotagged images instead of relying on structured datasets.
Uses Crowdsourced Validation (252 validators) to improve accuracy and reduce bias.
Detects small-scale environmental hazards that are undetectable by satellite-based AI models.
Incorporates Explainable AI (XAI) for greater transparency and journalistic accountability.
By bridging the gap between Computational Journalism, AI Fact-Checking, and Satellite AI, AIJIM represents a paradigm shift in AI-driven environmental journalism.
III-D AIJIM Workflow
AIJIM follows a structured four-step validation and reporting process:

### v5

# AIJIM: A Scalable Model for Real-Time AI in Environmental Journalism

@@ Abstract
This paper introduces AIJIM, the Artificial Intelligence Journalism Integration Model—a novel framework for integrating real-time AI into environmental journalism. AIJIM combines Vision Transformer-based hazard detection, crowdsourced validation with 252 validators, and automated reporting within a scalable, modular architecture. A dual-layer explainability approach ensures ethical transparency through fast CAM-based visual overlays and optional LIME-based box-level interpretations. Validated in a 2024 pilot on the island of Mallorca using the NamicGreen platform, AIJIM achieved 85.4% detection accuracy and 89.7% agreement with expert annotations, while reducing reporting latency by 40%. Unlike conventional approaches such as Data-Driven Journalism or AI Fact-Checking, AIJIM provides a transferable model for participatory, community-driven environmental reporting, advancing journalism, artificial intelligence, and sustainability in alignment with the UN Sustainable Development Goals and the EU AI Act.
1 Introduction
@@
Environmental journalism plays a critical role in raising awareness and shaping policy on issues such as climate change, pollution, and biodiversity loss, directly influencing public discourse and evidence-based decision-making [20, 25]. Yet, traditional methods suffer from slow data acquisition, costly verification, and limited scalability—particularly in under-monitored regions where urgent environmental threats (e.g., oil spills, illegal deforestation) require rapid response [12].
@@
Artificial Intelligence (AI)—including Natural Language Processing (NLP), Computer Vision (CV), and Machine Learning (ML)—has enhanced journalism through automation and large-scale analysis. However, existing approaches lack real-time adaptability in dynamic environmental contexts.
To address these gaps, we introduce the Artificial Intelligence Journalism Integration Model (AIJIM): a scalable, modular, and transferable model for integrating AI into real-time environmental journalism. AIJIM combines citizen-generated imagery, explainable AI (XAI), and crowdsourced validation to ensure precise detection, low latency, and transparent accountability. Its architecture—leveraging Vision Transformers and adaptable NLP pipelines—ensures future flexibility and cross-domain applicability.
In a 2024 pilot study on Mallorca, AIJIM demonstrated its effectiveness within the NamicGreen platform by analyzing 1,000 geotagged images, detecting 50 undocumented waste sites with 85.4% accuracy, and achieving 89.7% expert agreement—while reducing reporting latency by 40%.
Unlike existing models (e.g., Data-Driven Journalism, Computational Journalism), AIJIM bridges real-time detection, human-in-the-loop verification, and transparent reporting. This paper presents AIJIM’s conceptual foundation (Section 3), its practical implementation—demonstrated through detection overlays and heatmaps (Figs. 6–10)—and its implications for participatory, AI-driven journalism, while supporting the UN Sustainable Development Goals and compliance with the EU AI Act.
2 Related Work
@@
Existing research highlights the increasing integration of Artificial Intelligence (AI) in journalism, yet significant limitations persist. Traditional AI approaches, such as Data-Driven Journalism (DDJ), depend primarily on structured, retrospective datasets, rendering them ineffective for real-time environmental reporting [21, 35]. Similarly, Computational Journalism, which employs machine learning for complex pattern recognition, struggles with the dynamic and unstructured nature of live environmental data [13, 23]. In addition, AI Fact-Checking models, which focus on textual claim verification, lack capabilities for real-time visual or spatial hazard detection [37, 6].
@@
Building on these insights, it is essential to evaluate how existing AI-driven journalism frameworks address these challenges. This section identifies their methodological constraints and positions the Artificial Intelligence Journalism Integration Model (AIJIM) as a novel approach to integrating real-time automation, adaptability, and scalable journalistic workflows.
2.1 AI in Journalism
@@
AI technologies have significantly advanced journalism by enabling automation, rapid analysis, and streamlined verification processes. Natural Language Processing (NLP) has enhanced automated news generation from structured datasets, thereby improving journalistic efficiency [29, 35]. Additionally, developments in Computer Vision (CV) have facilitated rapid detection and classification of visual data, surpassing traditional manual methods in both speed and accuracy [34, 14]. Despite these advancements, current AI-driven methods often rely heavily on periodic updates from structured sources, such as satellite imagery, limiting their effectiveness in dynamic and rapidly evolving environmental scenarios. This gap underscores the need for AI solutions like AIJIM, which are capable of handling dynamic, real-time data streams efficiently.
2.2 Challenges in Environmental Journalism
@@
Environmental journalism frequently faces practical challenges, especially with regard to the timely acquisition and verification of reliable data. This is particularly evident in remote or under-monitored regions, complicating rapid responses to urgent environmental threats [25]. Conventional journalism methods—characterized by manual data collection and expert-driven verification—are resource-intensive, prone to inaccuracies, and hinder both scalability and timely intervention [21, 10]. Consequently, there is an urgent demand for innovative AI-driven solutions to improve the precision, timeliness, and overall credibility of environmental reporting.
2.3 Limitations of Existing AI Models
@@
Most AI models for environmental monitoring rely on satellite-based imagery and long-term trend analysis, which are effective for large-scale environmental assessments but insufficient for detecting localized, high-frequency hazards such as illegal dumping or micro-scale pollution [21, 35]. Additionally, existing AI solutions often fail to adjust adequately for contextual variability, resulting in misclassifications in underrepresented geographic regions [10]. Regulatory constraints—including GDPR compliance and data privacy risks—further limit the scalability of these models [14].
@@
2.4 Comparison of AIJIM with Existing Journalism Models
@@
Although several AI-based journalism models have been proposed, none fully meet the specific demands of real-time environmental crisis reporting. Data-Driven Journalism (DDJ), focused on historical structured datasets, lacks immediate adaptability [35]. Similarly, Computational Journalism, while proficient at analyzing extensive datasets, struggles with dynamic, real-time environmental scenarios [31]. Meanwhile, AI Fact-Checking excels primarily in textual verification but falls short in processing real-time visual and spatial data, which are critical for environmental monitoring [6].
@@
AIJIM addresses these limitations by uniquely integrating real-time hazard detection, adaptive community validation, and automated reporting into a coherent, modular framework. This combination ensures scalability, adaptability, and ethical AI governance, bridging the gap between automation and human oversight in environmental journalism. A detailed methodological comparison is presented in Section 3.
@@
Unlike purely technical frameworks or platform-specific toolkits, AIJIM is deliberately conceptualized as a theoretical model. Its goal is to offer a transferable structure that enables systematic analysis and design of AI-supported environmental journalism workflows. By integrating technological, ethical, and participatory dimensions, AIJIM contributes to ongoing efforts in theory-building at the intersection of journalism studies, human-centered AI, and digital environmental communication.
While model-agnostic explainability approaches like LIME and SHAP are well established in AI research, their practical application in real-time object detection remains limited due to performance constraints and architectural mismatch. In contrast, activation-based methods such as CAM offer high interpretability for convolutional models by visualizing spatial attention, and are therefore widely used in visual explanation pipelines. AIJIM combines both approaches by using CAM overlays as a fast default explainability layer and providing optional LIME-based per-box interpretation for expert validation.
3 The Artificial Intelligence Journalism Integration Model (AIJIM)
@@
The Artificial Intelligence Journalism Integration Model (AIJIM) advances beyond conventional AI-driven journalism by integrating real-time hazard detection, crowdsourced validation, and automated reporting into a scalable, modular, and adaptable framework. Unlike traditional models, AIJIM is designed for dynamic environmental journalism, enabling flexible integration of AI technologies tailored to specific reporting needs.
AIJIM’s modular architecture supports multiple AI components, ensuring adaptability. It integrates Vision Transformer-based DETR models, alternative CV and NLP architectures, and crowdsourced validation modules, allowing independent adjustment without system redesign [34]. This modularity enables AIJIM to address diverse environmental crises, from localized pollution to large-scale disasters, with rapid deployment across geographical regions.
AIJIM’s scalability supports high-volume data processing while maintaining accuracy and ethical responsibility. In the 2024 Mallorca pilot, AIJIM processed 1,000 citizen-generated images in hours, demonstrating efficiency without compromising quality [34]. Its adaptability integrates emerging technologies like satellite imagery or IoT sensors, enhancing capabilities in urban and remote areas.
Beyond detection, AIJIM enhances credibility through crowdsourced validation, achieving an 89.7% agreement rate with expert annotations in the Mallorca pilot [34]. For reporting, AIJIM employs adaptable NLP models, reducing latency by 40% [29]. By converging automation and human oversight, AIJIM delivers scalable, accurate, and ethically responsible journalism, applicable across investigative and large-scale monitoring contexts.
3.1 Methodology
@@
AIJIM’s effectiveness was evaluated using a Sequential Explanatory Mixed-Methods Approach [8], combining quantitative and qualitative analyses:
Quantitative Analysis: The 2024 Mallorca pilot tested AIJIM’s real-time capabilities, analyzing 1,000 geotagged citizen images to detect 50 undocumented waste sites [34, 27]. The NamicGreen platform served as a proof-of-concept, with the DETR-based Vision Transformer enabling robust detection, though AIJIM supports alternative models.
@@
Qualitative Analysis: Twelve expert interviews with journalism, AI, and regulatory professionals informed the evaluation. Structured Qualitative Content Analysis (QCA) confirmed thematic dimensions with a Krippendorff’s Alpha of 85%, ensuring reliability [26].
3.2 Technology Adaptability
@@
AIJIM’s architecture supports diverse AI models and hardware environments. Its modular structure enables integration of alternative object detection and NLP models based on deployment needs, ensuring scalability and flexibility. As a technology-agnostic model, AIJIM adopts future AI advancements without structural changes, optimizing performance across journalistic contexts [34].
3.3 Comparison with Existing AI-Based Journalism Models
AIJIM surpasses traditional AI-driven journalism models—such as Data-Driven Journalism (DDJ), Computational Journalism, AI Fact-Checking, and Satellite AI—by enabling real-time, multi-modal environmental reporting. Unlike DDJ, which relies on retrospective datasets, AIJIM processes unstructured, citizen-generated data instantly, providing immediate environmental insights. AIJIM’s modular design allows the integration of various AI models for enhanced adaptability, addressing limitations in existing frameworks. A detailed comparison across key dimensions is presented in Table 1 and Table 2 in Appendix A.
AI Fact-Checking verifies textual claims but lacks visual hazard detection [11]. AIJIM integrates multi-modal analysis with crowdsourcing.
As shown in Fig. 1, AIJIM surpasses existing models in real-time processing, public engagement (252 validators), and small-scale hazard detection, supported by Explainable AI (XAI) for transparency [32]. The 2024 Mallorca pilot validated these strengths, achieving 40% latency reduction and 85.4% accuracy [34], establishing a new benchmark for scalable, participatory journalism.
3.4 AIJIM Workflow
AIJIM follows a six-step validation and reporting process, as illustrated in Fig. 2: